If 450000 = 3 to the power of x, then how do you solve for x without using the graphing features on a graphing calculator?
Answer by stanbon(75887)

If 450000 = 3 to the power of x, then how do you solve for x without using the graphing features on a graphing calculator?
-----
450000 = 3^x
----
Use the log features.
x(log3) = log(450000)
x = log(450000)/log3
x = 11.8486..
==================
Cheers,
Stan H.